ABC is a triangle in which AB=AC. A circle through B touches AC at D intersects AB at P.If D is the mid point of AC, prove that AB=4AP.
Thank You

Answer by ly12603(44)

step 1: 4(x-3)=3(x-4)
step 2: 4x-12=3x-12 <---distribute
step 3: -3x -3x <---solve like a normal equation
_______________
answer: x-12=-12
step 4: +12 +12
_________________
answer: x=0 <---try checking it it works :)